# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v2

@@ III.3 The upgoing solution
$C_{\rm up}$ $=$ ${(2M)^{3}e^{-\partial_{\kappa_{1}}{\cal F}_{\rm inst}}\over x^{3+\frac{x}{2}}B_{++}}\left(1-{\textstyle\frac{B_{-+}B_{+-}}{B_{++}B_{--}}}\right)^{-1}=$ (67)
$=\frac{(2M)^{3}}{x^{3+\frac{x}{2}}}\frac{\Gamma(2-\ell)}{\Gamma(-2\ell)}+\ldots$
@@
IV Tests against MST and numerical simulations
In this Section we test our hypergeometric PM representation against the MST method [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and numerical simulations. When we restrict ourselves to circular orbits [^2] As it is well known in this case the PM and PN expansion are just one expansion., we work at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the probe particle. We use the package Teukolsky of the Black Hole Perturbation Toolkit (BHPT) [56] both for the computations using the MST method and for numerical integration of the differential equation. The observables we consider are the $(\ell,m)$ harmonic modes of incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system. The results for the energy flux are also compared against the 22PN results in [20].
The numerical computations and those with the MST method are performed setting in the BHPT Mathematica files: AccuracyGoal = 1000, PrecisionGoal = 70, WorkingPrecision = 80, SetPrecision=80. With this choice of parameters both numerical and MST results agree to a 15 digits precision. We restrict our comparisons to circular orbits for the aforementioned observables, so $r$ should be taken bigger than the innermost stable circular orbit (ISCO) radius $r_{\text{ISCO}}=6M$. Consequently the domain of variability of the tangential velocity is $v\in[0,\sqrt{\frac{M}{r}}=\frac{1}{\sqrt{6}}]$.
IV.1 The incoming and upgoing solutions
In this subsection we collect all the ingredients needed in the computations of the incoming, upgoing waveforms at order k=30 in the hypergeometric PM representation. We write333 Here we rewrite the second terms in the right hand sides of (62), (66), as the images under $a\to-a$ of the first terms, using the transformation rules $R_{-}\to R_{+}$, $\gamma\to\gamma^{-1}$, $F_{-\alpha^{\prime}}\leftrightarrow F_{+\alpha^{\prime}}$ and $B_{-\alpha^{\prime}}\leftrightarrow B_{+\alpha^{\prime}}$.
@@ IV.1 The incoming and upgoing solutions
with
$R_{\alpha}(y)$ $=$ $e^{-{y\over 2}}(1-{\textstyle\frac{x}{y}})^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}\left[P_{0}(y)\,H^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)\,yH^{0}_{\alpha}{}^{\prime}(y)\right]$
@@
$C_{\rm up}$ $=$ ${(2M)^{3}\over x^{3+\frac{x}{2}}\,\gamma_{\kappa}}{\Gamma({\textstyle\frac{5}{2}}-a+{\textstyle\frac{x}{2}})\over\Gamma(1-2a)}\left(1-e^{-2\pi{\rm i}a}\,{\textstyle\frac{\cos{\textstyle\frac{\pi}{2}}(x+2a)}{\cos{\textstyle\frac{\pi}{2}}(x-2a)}}\right)^{-1}$
$\gamma_{\kappa}$ $=$ $1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}$
$C_{\rm in}$ $=$ $x^{x\over 2}e^{-\frac{{\rm i}\pi}{2}(5+2a+x)}{\Gamma(-{\textstyle\frac{3}{2}}+a-{\textstyle\frac{x}{2}})\over\Gamma(1+2a)}\left(1-\gamma_{0}\,\gamma\right)^{-1}$ (70)
and [^4] We notice that $\gamma_{\rm tot}=e^{-\partial_{a}({\cal F}_{\rm tree}+{\cal F}_{\rm one-loop}+{\cal F}_{\rm inst})}$ collects the tree level, one-loop and instanton contributions to the prepotential of the underlying $SU(2)$ gauge theory with three fundamental flavours.
$\gamma_{\rm tot}=\underbrace{e^{-2\pi{\rm i}a}\,\frac{\Gamma({-}2a)^{2}\Gamma\left({-}\frac{3}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{+}a{-}\frac{x}{2}\right)}{\Gamma(2a)^{2}\Gamma\left(-\frac{3}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{-}a{-}\frac{x}{2}\right)}}_{\gamma_{0}=\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}\underbrace{x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}}}_{\gamma}\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\cal F}_{\rm inst}=-\sum_{i=1}^{\infty}{u_{i}\over i}x^{i}$ (71)
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\,,\,\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (72)
@@
$A_{1}$ $=$ $-8\left(\Delta_{ij}+1\right)\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{A}_{1}=4\left(4a^{2}-(x+5)^{2}\right)\Delta_{i,j}$
$A_{3}$ $=$ $8\Delta_{ij}\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{A}_{3}=4\left(4a^{2}-(x+5)^{2}\right)\left(-\Delta_{i,j}+x-2\right)$
@@
$B_{1}$ $=$ $-16\left(\Delta_{ij}+1\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{B}_{1}=8\Delta_{ij}\left(\Delta_{ij}+x+5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{B}_{3}=8\left(-\Delta_{i,j}^{2}-7\Delta_{ij}+x^{2}+3x-10\right)$
$B_{2}$ $=$ $4\left(-2(x-1)\Delta_{i,j}+4a^{2}+x^{2}-2x-3\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ B_{3}=16\Delta_{ij}$
@@
$C_{1}$ $=$ $8\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{C}_{1}=16a^{2}-4(x+5)^{2}\leavevmode\nobreak\ ,\leavevmode\nobreak\ C_{2}=16\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{C}_{2}=8\left(\Delta_{ij}+x+5\right)$ (74)
@@
Finally one has to compute $a(\ell)$. Given $u_{i}(a)$ by (73), one can invert the series (12) to compute $a(u)$ and then set (from (41))
@@
Inverting the series at high PM order can be time consuming for a Mathematica code. There is an alternative, more direct way to find the function $a(u)$ based on the connection between this quantity and the quantum SW period of a $SU(2)$ theory with three fundamental flavours. The quantum SW period $a$ is known to satisfy the continuous fraction equation [57]
@@ IV.2 The luminosity
the energy and angular momentum of the light particle. Finally, the luminosity, i.e. the amount of energy per unit time emitted by the system towards infinity, is computed by the formula (see [16] and references therein)
@@ IV.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHPT [56]. In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
$k$ $v=0.1$ $v=0.3$ $v=0.4$
5 0.961483706187661 0.841473688330721 0.887926055076
510 0.961483705502195 0.841176615502628 0.880582026053
15 0.961483705502172 0.841176157785812 0.880686334361
20 0.961483705502172 0.841176171311697 0.880705358744
25 0.961483705502172 0.841176171325155 0.880705412455
30 0.961483705502172 0.841176171325065 0.880705411714
MST/NUM 0.961483705502277 0.841176171325069 0.880705411627
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in [20].
@@ V The Heun equation
$G^{\prime\prime}(z)+\left({\gamma_{0}\over z}+{\gamma_{1}\over z-1}+{\gamma_{2}\over z-x}\right)G^{\prime}(z)+{-\sigma+z\beta_{0}\beta_{1}\over z(z-1)(z-x)}G(z)=0$ (86)
The parameters satisfy the linear relation $\gamma_{0}+\gamma_{1}+\gamma_{2}=1+\beta_{0}+\beta_{1}$. The characteristic exponents around $z\approx x$ are $(0,1-\gamma_{2})$ and at infinity are $(\beta_{0},\beta_{1})$. The solutions of the Heun equation can be written as linear combinations of
${\rm Heun}_{-}(z)$ $=$ ${\rm Heun}\left[x,\sigma,\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},z\right]$ (87)
${\rm Heun}_{+}(z)$ $=$ $z^{1{-}\gamma_{0}}{\rm Heun}\left[x,\sigma{+}\left(1{-}\gamma_{0}\right)\left(\gamma_{2}{+}\gamma_{1}x\right),1{+}\beta_{0}{-}\gamma_{0},1{+}\beta_{1}{-}\gamma_{0},2{-}\gamma_{0},\gamma_{1},z\right]$
where ${\rm Heun}\left[x,\sigma,\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},z\right]$ is the Heun function which is normalized to one at $z=0$ so that
${\rm Heun}_{-}(z)=1+\ldots\,,\,{\rm Heun}_{+}(z)=z^{1-\gamma_{0}}(1+\ldots)$ (88)
The dots stand for corrections in $z$. The expansion coefficients of ${\rm Heun}_{\alpha}(z)$ can be obtained recursively, solving the equation order by order in $z$. As before, we find it convenient to parameterize the characteristic exponents $\beta_{i},\sigma$ of the Heun function in terms of the variables $\kappa_{i}$, $\widehat{u}$, defined as
$\beta_{0}$ $=$ $\kappa_{1}{+}\kappa_{3}\,,\,\beta_{1}=\kappa_{1}{+}\kappa_{2}\,,\,\sigma=\kappa_{1}^{2}-\widehat{u},$
$\gamma_{1}$ $=$ $\kappa_{2}{+}\kappa_{3}\,,\,\gamma_{2}=1{-}\gamma_{0}+2\kappa_{1}$ (89)
@@
$G(z)=\sum_{\alpha}B_{\alpha}\,{\rm Heun}_{\alpha}(z)$ (90)
@@
$G(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1{-}z)^{1-\kappa_{2}-\kappa_{3}}(1+\ldots)$
$G(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{1-\gamma_{0}}(1+\ldots)$ (91)

### v3

@@ 3.3 The upgoing solution
$C_{\rm up}$ $=$ ${(2M)^{3}e^{\partial_{m_{3}}{\cal F}_{\rm inst}}\over x^{3+\frac{x}{2}}B_{++}}\left(1-{\textstyle\frac{B_{-+}B_{+-}}{B_{++}B_{--}}}\right)^{-1}=\frac{(2M)^{3}}{x^{3+\frac{x}{2}}}\frac{\Gamma(2-\ell)}{\Gamma(-2\ell)}+\ldots$ (71)
@@
4 Wave form and energy flux at 30PN
In this Section we compute the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN, we mean that the polynomials coefficients $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
Since for circular orbits $v^{2}=M/r$, and $r$ should be taken bigger than the innermost stable circular orbit (ISCO) radius $r_{\text{ISCO}}=6M$, the domain of variability of the tangential velocity is $v\in[0,{\textstyle\frac{1}{\sqrt{6}}}]$. We find that the first 12 digits of 30PN results are stable for any choice $v$ in this domain, so we will always display this number of digits. The numerical computations and those with the MST method are performed setting in the BHTK Mathematica files: AccuracyGoal = 1000, PrecisionGoal = 70, WorkingPrecision = 80, SetPrecision=80. With this choice of parameters both numerical and MST results agree to a 12 digits precision. In our tables these results will be denoted as BHTK.
4.1 The incoming and upgoing solutions
In this subsection we collect all the ingredients needed in the computations of the incoming, upgoing waveforms at order k=30 in the hypergeometric PM representation. We write555 Here we rewrite the second terms in the right hand sides of (66), (70), as the images under $a\to-a$ of the first terms, using the transformation rules $R_{-}\to R_{+}$, $\gamma\to\gamma^{-1}$, $F_{-\alpha^{\prime}}\leftrightarrow F_{+\alpha^{\prime}}$ and $B_{-\alpha^{\prime}}\leftrightarrow B_{+\alpha^{\prime}}$.
@@ 4.1 The incoming and upgoing solutions
with [^6] We notice that $\gamma_{\rm tot}=\gamma{\textstyle\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}=e^{-2\pi{\rm i}a_{D}}=e^{-\partial_{a}({\cal F}_{\rm tree}+{\cal F}_{\rm one-loop}+{\cal F}_{\rm inst})}$ collects the tree level, one-loop and instanton contributions to the dual SW period $a_{D}$.
$R_{\alpha}(y)$ $=$ $e^{-{y\over 2}}\left(1-{\textstyle\frac{x}{y}}\right)^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}\left[P_{0}(y)\,H^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)\,yH^{0}_{\alpha}{}^{\prime}(y)\right]$
@@
$C_{\rm in}$ $=$ $x^{x\over 2}e^{-\frac{{\rm i}\pi}{2}(5+2a+x)}{\Gamma(-{\textstyle\frac{3}{2}}+a-{\textstyle\frac{x}{2}})\over\Gamma(1+2a)}\left(1-\gamma_{\rm tot}\right)^{-1}$
$C_{\rm up}$ $=$ ${(2M)^{3}\over x^{3+\frac{x}{2}}\,\gamma_{m}}{\Gamma({\textstyle\frac{5}{2}}-a+{\textstyle\frac{x}{2}})\over\Gamma(1-2a)}\left(1-e^{-2\pi{\rm i}a}\,{\textstyle\frac{\cos{\textstyle\frac{\pi}{2}}(x+2a)}{\cos{\textstyle\frac{\pi}{2}}(x-2a)}}\right)^{-1}$
$\gamma_{m}$ $=$ $1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}\quad,\quad\gamma=x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}}$
$\gamma_{\rm tot}$ $=$ $\gamma\,e^{-2\pi{\rm i}a}\,\frac{\Gamma({-}2a)^{2}\Gamma\left({-}\frac{3}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{+}a{-}\frac{x}{2}\right)}{\Gamma(2a)^{2}\Gamma\left(-\frac{3}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{-}a{-}\frac{x}{2}\right)}$ (74)
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\quad,\quad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (75)
@@
$A_{1}$ $=$ $-8\left(\Delta_{ij}+1\right)\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{1}=4\left(4a^{2}-(x+5)^{2}\right)\Delta_{i,j}$
$A_{3}$ $=$ $8\Delta_{ij}\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{3}=4\left(4a^{2}-(x+5)^{2}\right)\left(-\Delta_{i,j}+x-2\right)$
@@
$B_{1}$ $=$ $-16\left(\Delta_{ij}+1\right)~{},~{}\widehat{B}_{1}=8\Delta_{ij}\left(\Delta_{ij}+x+5\right)~{},~{}\widehat{B}_{3}=8\left(-\Delta_{i,j}^{2}-7\Delta_{ij}+x^{2}+3x-10\right)$
$B_{2}$ $=$ $4\left(-2(x-1)\Delta_{i,j}+4a^{2}+x^{2}-2x-3\right)~{},~{}B_{3}=16\Delta_{ij}$
@@
$C_{1}$ $=$ $8\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{C}_{1}=16a^{2}-4(x+5)^{2}~{},~{}C_{2}=16~{},~{}\widehat{C}_{2}=8\left(\Delta_{ij}+x+5\right)$ (77)
@@
Finally one has to compute $a(\ell)$. Given $u_{i}(a)$ by (76), one can invert the series (2.1) to compute $a(u)$ and then set (from (45))
@@
Inverting the series at high PM order can be time consuming for a Mathematica code. There is an alternative, more direct way to find the function $a(u)$ based on the connection between this quantity and the quantum SW period of a SU(2) theory with three fundamental flavours. The quantum SW period $a$ is known to satisfy the continuous fraction equation Poghosyan:2020zzg
@@ 4.2 The luminosity
the energy and angular momentum of the light particle. Finally, the luminosity, i.e. the amount of energy per unit time emitted by the system towards infinity, is computed by the formula (see Mino:1997bx and references therein)
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHTK BHPToolkit . In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in Fujita:2012cm .
Method $v=0.1$ $v=0.3$ $v=0.4$
22PN 0.961483705502 0.841176171326 0.880705608981
HYP 0.961483705502 0.841176171325 0.880705411714
BHTK 0.961483705502 0.841176171325 0.880705411627
@@ 5 The Heun equation
$G_{1}^{\prime\prime}(z)+\left({\beta_{0}\over z}+{\beta_{1}\over z-1}+{\beta_{2}\over z-\lambda}\right)G_{1}^{\prime}(z)+{z\alpha_{0}\alpha_{1}-\sigma\over z(z-1)(z-\lambda)}G_{1}(z)=0$ (89)
with $\beta_{2}=1+\alpha_{0}+\alpha_{1}-\beta_{0}-\beta_{1}$. The solutions of the Heun equation can be written as linear combinations of
${\rm Heun}_{-}(z)$ $=$ ${\rm Heun}\left[\lambda,\sigma,\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},z\right]$ (90)
${\rm Heun}_{+}(z)$ $=$ $z^{1{-}\beta_{0}}{\rm Heun}\left[\lambda,\sigma{+}\left(1{-}\beta_{0}\right)\left(\beta_{2}{+}\lambda\beta_{1}\right),1{+}\alpha_{0}{-}\beta_{0},1{+}\alpha_{1}{-}\beta_{0},2{-}\beta_{0},\beta_{1},z\right]$
where ${\rm Heun}\left[\lambda,\sigma,\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},z\right]$ is the Heun function which is normalized to one at $z=0$, so that
${\rm Heun}_{-}(z)=1+\ldots\,,\,{\rm Heun}_{+}(z)=z^{1-\beta_{0}}(1+\ldots)$ (91)
The dots stand for corrections in $z$. The expansion coefficients of ${\rm Heun}_{\alpha}(z)$ can be obtained recursively, solving the equation order by order in $z$. The Heun differential equation can be related to the quantum SW curve of a ${\cal N}=2$ supersymmetric gauge theory with four fundamental hypermultiplets. To this aim, we write the six Heun parameters in terms of 4 masses $m_{i}$, a gauge coupling $x$ and a Coulomb branch parameter $u$, defined via the identifications777 Equation (89) can be alternatively written in the normal form (5) with $G_{1}(z)=(1-z)^{-\frac{\beta_{1}}{2}}z^{-\frac{\beta_{0}}{2}}(z-\lambda)^{-\frac{\beta_{2}}{2}}\,\Psi(z)$ (92) and $Q_{\rm gauge}$ $=$ $\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}+m_{4}}{2}}\right){}^{2}}{(z-x)^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}-m_{4}}{2}}\right){}^{2}}{z^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{1}+m_{2}}{2}}\right){}^{2}}{(z-1)^{2}}+\frac{m_{3}^{2}+m_{4}^{2}+2m_{1}m_{2}-1}{2(z-1)z}$ (93) $+\frac{(1-x)\left(u+{\textstyle\frac{1}{4}}-\frac{m_{3}^{2}}{2}-\frac{m_{4}^{2}}{2}\right)-{\textstyle\frac{x}{2}}\left(1{-}m_{1}{-}m_{2}\right)\left(1{-}m_{3}{-}m_{4}\right)}{(z-1)z(z-x)}$
$\alpha_{0}$ $=$ $1{-}m_{1}{-}m_{3}\quad,\quad\alpha_{1}=1{-}m_{2}{-}m_{3}\quad,\quad\lambda=x$ (94)
$\beta_{0}$ $=$ $1{-}m_{3}{+}m_{4}\quad,\quad\beta_{1}=1{-}m_{1}{-}m_{2}\quad,\quad\beta_{2}=1-m_{3}-m_{4}$
$\sigma$ $=$ $u(x-1)+\left({\textstyle\frac{1}{2}}-m_{3}\right){}^{2}+x\left({\textstyle\frac{3}{4}}{-}m_{1}{-}m_{2}{-}m_{3}{+}m_{1}m_{2}{+}m_{1}m_{3}{+}m_{2}m_{3}\right)$
@@
$G_{1}(z)=\sum_{\alpha}B_{\alpha}\,{\rm Heun}_{\alpha}(z)$ (95)
@@
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1{-}z)^{m_{1}+m_{2}}(1+\ldots)$
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{m_{3}-m_{4}}(1+\ldots)$ (96)